OpenStudy (anonymous):

Question about how to re-arrange a complicated formula. re-arranging s=vt - 0.5*at^2 to make T the subject @Engineering

OpenStudy (anonymous):

at^2 -2vt +2s = 0

OpenStudy (anonymous):

I just don't understand how to get there

OpenStudy (anonymous):

s=vt - 0.5*at^2 s= vt -1/2(at²) -at²/2+vt-s=0 (-a*t²)/2+v*t-s=0

OpenStudy (anonymous):

thnak you very much do you have any general tips on re-arranging formulae? i know you do the opposite to what you have, but what about the order?

OpenStudy (anonymous):

Just try to put it under the following form: \[ax^{2} +bx + c =0\]or \[ax^{n} +bx^{n-1}+ ... + yx+z=0\]and review the steps a followed in my first post.

OpenStudy (anonymous):

thank you very much